The ionic conductivity of Ba++ plus 2 and cl- at infinite dilution are 127 and 76 ohm inverse centimetre square equivalent inverse respectively. The equivalent conductivity of bacl2 at Infinity dilution:

Grade:12th pass

2 Answers

Adarsh
6 years ago
Dear Akash,
At infinite dilution each ion contributes towards the equivalent conductance of the electrolyte .. The value of equivalent conductance at infinite dilution for any electrolyte is the sum of contribution of its constituent ions..
so equivalent conductance = 127 + 76 = 203.
Thanks and Regards
Adarsh
Ak
4 years ago
0.5(127)+76 = 139.5. This is the equivalent conductance of bacl2 at infinite dilution. Equivalent conductance at infinite dilution of BA 2 bacl2 is half of that of Ba 2+ and one by one of that of Cl-